### v1

# TimeMIL: Advancing Multivariate Time Series Classification via a Time-aware Multiple Instance Learning

Abstract
Deep neural networks, including transformers and convolutional neural networks, have significantly improved multivariate time series classification (MTSC). However, these methods often rely on supervised learning, which does not fully account for the sparsity and locality of patterns in time series data (e.g., diseases-related anomalous points in ECG). To address this challenge, we formally reformulate MTSC as a weakly supervised problem, introducing a novel multiple-instance learning (MIL) framework for better localization of patterns of interest and modeling time dependencies within time series. Our novel approach, TimeMIL, formulates the temporal correlation and ordering within a time-aware MIL pooling, leveraging a tokenized transformer with a specialized learnable wavelet positional token. The proposed method surpassed 26 recent state-of-the-art methods, underscoring the effectiveness of the weakly supervised TimeMIL in MTSC. The code will be available at https://github.com/xiwenc1/TimeMIL.
Multivariate Time Series Classification, Weakly-supervised Learning
1 Introduction
Time series data mining has witnessed considerable growth in the last decade with numerous applications in classification (Ismail Fawaz et al., 2019), forecasting (Lim & Zohren, 2021), and anomaly detection (Malhotra et al., 2015). Particularly, multivariate time series classification (MTSC), which aims to assign labels to time sequences, is challenging but crucial in most real scenarios, such as healthcare (Vrba & Robinson, 2001; Tang et al., 2023), human action recognition (Shokoohi-Yekta et al., 2017; Amaral et al., 2022), audio signal processing (Ruiz et al., 2021), Internet of Things (Bakirtzis et al., 2022), and semantic communication (Zhao et al., 2023).
Figure 1: (a): The decision boundary of fully supervised methods is determined by assigning a label to each time series. (b): TimeMIL makes decisions by discriminating positive and negative instances in time series, where each time point is an instance, and its label is typically not available in reality.
Recently, deep neural networks have achieved state-of-the-art performance in various time series tasks compared to traditional methods (Seto et al., 2015; Schäfer & Leser, 2017; Li et al., 2023a; Tang et al., 2022; donghao & wang xue, 2024; Li et al., 2024a). Their popularity and success in time series modeling can be attributed to their automatic feature extraction in conjunction with inductive biases. To this end, the time series task is typically formulated as a fully supervised learning task by employing a wide variety of architectures, such as recurrent neural networks (RNN) (Franceschi et al., 2019; Lai et al., 2018b), long short-term memory (LSTM), (Karim et al., 2019; Tang et al., 2022; Karim et al., 2019), convolution neural networks (CNN)  (Zhang et al., 2020; Ismail Fawaz et al., 2020; Wu et al., 2023; Tang et al., 2022; donghao & wang xue, 2024), and transformers (Zhou et al., 2022; Zhang & Yan, 2023; Nie et al., 2023; Li et al., 2024b, c). Although these methods vary in their architectural biases for modeling time series, they share a common underlying strategy: dividing a time series into a set of time points and then modeling the local and global dependencies among them, such as changes over time and the presence of multiple periodic patterns. Importantly, patterns of interest in time series are typically sparse and localized (Lin et al., 2012; Jiang et al., 2011; Fulcher, 2018; Cheng et al., 2009), while the most discriminative time points within a time series are typically unknown due to their laborious annotation. This poses a significant challenge for fully supervised learning in accurately determining the decision boundary (see Fig. 1(a)). Instead, given the inherent properties of time series, we formulate the MTSC tasks as a weakly supervised learning paradigm (see Fig. 1(b)).
Multiple Instance Learning (MIL) is a weakly supervised learning method that assigns a label to a collection of instances, known as a bag. This makes MIL a natural choice for the MTSC task by collectively treating each time point as an instance and an entire time series as a bag. Early attempts of MIL in time series relied on hand-crafted features and classic MIL models (Stikic et al., 2011; Guan et al., 2016). In contrast, modern MILs which use deep neural networks to extract the feature automatically and consistently exhibit superior performance compared to the classic MILs with handcrafted features (Wang et al., 2018; Ilse et al., 2018; Early et al., 2024). However, standard MILs may fail to capture correlations between instances, since standard MILs assume the independence and identical distribution of instances with a permutation-invariant property (Ilse et al., 2018). In contrast, MTSC data typically exhibits temporal correlations and ordering dependencies, posing significant challenges for directly translating MIL into MTSC.
This paper introduces a generic MIL framework for time series, termed TimeMIL. We address several limitations when using standard MIL methods, such as their failure to model the permutation information and temporal correlation among instances. We explore their necessity from an information-theoretic perspective, which suggests that modeling the permutation information and temporal correlation can lower the uncertainty of classification systems. To this end, we propose a time-aware MIL pooling, leveraging the self-attention mechanism and a novel learnable wavelet positional encoding, where the former is used to capture the temporal correlation between instances, and the latter is used to characterize time ordering information.
Contributions: (i) To the best of our knowledge, we are the first to formally formulate a generic MIL framework for multivariate and multi-class time series classification from an information-theoretic perspective. (ii) We propose a Time-aware MIL pooling based on a tokenized transformer and a novel learnable wavelet positional encoding to model complex patterns within time series. The proposed method outperforms 26 recent state-of-the-art methods in 28 datasets and offers inherent interpretability.
2 Related Works
Multivariate Time Series Classification. The recent DL methods specifically designed for MTSC can roughly be divided into two categories: (i) CNN/LSTM Hybrid architecture (Karim et al., 2019; Zhang et al., 2020), where LSTM is often used to capture sequential dependencies and CNN is used to capture the inter-variate correlations. (ii) Purely CNN architecture (Ismail Fawaz et al., 2020; Li et al., 2021b; Tang et al., 2022), where long-term dependencies, short-term dependencies, and cross-channel dependencies are claimed to be captured by multiple kernels with varying kernel sizes.
Recently, the authors of (Wu et al., 2023; donghao & wang xue, 2024) offered a model called General Time Series Analysis Framework, to be used in multiple mainstream tasks, including classification, imputation, short-term forecasting, long-term forecasting, and anomaly detection, with simple modifications. Transformer-based models (Zhou et al., 2022; Zhang & Yan, 2023; Nie et al., 2023) and MLP-based models (Zeng et al., 2023; Zhang et al., 2022b; Li et al., 2023c, b) have also been developed and improved over the last few years for this purpose due to their excellent scaling behaviors. Nonetheless, they did not yet fully replace CNN-based models, which continue to exhibit impressive performance (Liu et al., 2022; Wang et al., 2023; donghao & wang xue, 2024).
Although the aforementioned methods have witnessed extensive use in time series analysis applications, as discussed in the introduction, these methods are rooted in supervised learning and often focus on optimizing their architectural designs (e.g., CNN and transformer), which still cannot solve the essential issue that they are challenging to determine the accurate decision boundary. In contrast, our work introduces TimeMIL, which provides a novel perspective to describe the decision boundary of time series in a weakly supervised view. In addition, the proposed TimeMIL is inherently interpretable.
Figure 2: The proposed framework of TimeMIL for time series classification with enhanced interpretability: (i) a feature extractor to obtain instance-level feature embeddings, (ii) a MIL pooling to aggregate instance embeddings to a bag-level feature, embedding, and (iii) a bag-level classifier to map bag-level feature to a label prediction. Each time point is treated as an instance and the time series as a bag. Time ordering information and instance correlation are captured by taking the mutual benefit of WPE and MHSA in our TimeMIL pooling (highlighted in green).
Multiple Instance Learning. MIL, a weakly supervised method, is widely used for histological image classification due to its advantage in localizing tumors within gigapixel images (Ilse et al., 2018; Li et al., 2021a; Zhang et al., 2022a). However, the application of MIL to time series data has rarely been explored. An early exploration by (Stikic et al., 2011) applied classic multi-instance SVM (Andrews et al., 2002) to wearable sensor data. However, this method fails to model sequential dependencies among time points (instances). To address this limitation, (Guan et al., 2016) proposes incorporating an autoregressive hidden Markov model into the MIL framework to model the dependencies between instances. However, both methods rely on FFT and hand-crafted statistical features. Alternatively, modern MILs typically employ deep neural networks for automatic feature extraction and aggregation (Wang et al., 2018), achieving superior performance over conventional instance-based MIL, like the one used in (Zhu et al., 2021) for MTSC. Most notably, the attention-based MIL (ABMIL) is proposed by (Ilse et al., 2018), which makes MIL inherently interpretable by weighting each instance according to its importance. Since its invention, ABMIL has emerged as the standard paradigm for modern MIL applications (Li et al., 2021a; Zhu et al., 2023; Shao et al., 2021; Zhang et al., 2022a; Qiu et al., 2023; Xiang & Zhang, 2023). Following this line of work, a related but concurrent work (Early et al., 2024) attempted to apply ABMIL and its variants to time series. However, ABMIL operates under the assumption that instances are independent and identically distributed, which inherently limits the modeling of the temporal dependencies among instances in time series data.
While being related to methods presented in  (Guan et al., 2016; Early et al., 2024), our method differs from them in the following ways. (i) Unlike the autoregressive hidden Markov used in (Guan et al., 2016), which struggles with long-range and complex dependencies, the proposed method employs self-attention to model the instance dependencies regardless of their distance, offering inherent interpretability. (ii) Authors of (Early et al., 2024) directly applied MIL to the time series classification problem to obtain interpretability without providing a theoretical justification of how time series classification can be framed as a MIL problem. Especially for the multi-class cases, they violated the MIL assumption that a bag is positive as long as a positive instance is present. Additionally, their proposed method also falls into the category of ABMIL, which does not naturally model the temporal correlation and ordering of time points within a time series. In contrast, we re-frame the MIL for more complex multi-class MTSC tasks. Specifically, we show how to effectively tackle multi-class problems in the context of the binary MIL paradigm (Section 3.2), clarifying and addressing the limitations of standard (AB)MIL techniques used for time series analyses (Section 3.3).
3 Method
In this section, we introduce three key components for applying the proposed TimeMIL to MTSC. First, we formulate the MTSC as a MIL problem in section 3.1 and 3.2. Second, we introduce a time-aware MIL pooling to capture the temporal ordering in time series through a wavelet-augmented transformer (Section 3.3). The entire framework of the proposed TimeMIL is depicted in Fig. 2. Third, we introduce interpretable pooling to assess the importance of instance, as discussed in section 3.4. A summary of the proposed framework is presented in Algorithm 1.
3.1 Problem Formulation
Multivariate Time Series data is typically presented as $\{\boldsymbol{X}_{1},\cdots,\boldsymbol{X}_{n}\}$, where $\boldsymbol{X}_{i}=\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}$ is a time series with $T$ time points, with each time point $\boldsymbol{x}_{i}^{t}\in\mathbb{R}^{d}$ being a $d$-dimensional vector. It is noteworthy that the time points are shift-variant and ordered. Similar to the standard classification task, the goal of MTSC is to learn a direct mapping from feature space $\mathcal{X}$ to label space $\mathcal{Y}$ using the training data set $\{(\boldsymbol{X}_{1},y_{1}),\cdots,(\boldsymbol{X}_{n},y_{n})\}$, where $y_{i}$ is the label for a single time series.
3.2 MTSC as a MIL problem
Binary MTSC. Without violating the MIL assumption, we take binary MTSC as a starting point in the following derivations, then extend it to the multi-class scheme. The goal of a binary MIL is to assign a label to a bag of instances. The MTSC can naturally be formulated as a MIL problem by treating each time series as a bag, with each time slot being an instance. Formally, the binary MTSC under the MIL formulation is defined as
$y_{i}=\left\{\begin{array}[]{l}0,\text{ iff }\sum_{t=1}^{T}y_{i}^{t}=0,\ y_{i}^{t}\in\{0,1\}\\ 1,\text{ otherwise, }\end{array}\right.$ (1)
where $y_{i}^{t}$ denotes the label for each time point indicating if an event of interest has happened at time point $\boldsymbol{x}_{i}^{t}$. The $\{y_{i}^{1},\cdots,y_{i}^{t}\}$ are also known as the instance-level labels in the context of MIL, which are unknown in most scenarios. Eq. 1 implies that a time series (bag) $\boldsymbol{X}_{i}$ is labeled as positive if and only if any of its instance labels is positive, negative otherwise.
The bag-level prediction $\hat{y}_{i}$ of a MIL is given as a score function $S:\mathcal{X}\rightarrow\mathbb{R}$ (Ilse et al., 2018):
$\hat{y}_{i}=S(\boldsymbol{X}_{i}),$ (2)
where the outcome of a score function is a probability.
Theorem 1.
(Ilse et al., 2018; Shao et al., 2021) Suppose the score function $S$ is a $(\delta_{\varepsilon},\varepsilon)$-continuous symmetric function w.r.t Hausdorff distance $d_{H}(\cdot,\cdot)$, i.e. $\forall d_{H}(\boldsymbol{X}_{i},\boldsymbol{X}_{j})<\delta_{\varepsilon}$, we have $|S(\boldsymbol{X}_{i})-S(\boldsymbol{X}_{j})|<\varepsilon$, for $\forall\varepsilon>0$. For any invertible map $\sigma:\mathcal{X}\rightarrow\mathbb{R}^{d}$, $S$ can be approximated by certain continuous functions $g$ and $f$:
$|S(\boldsymbol{X}_{i})-g(\sigma\{f(\boldsymbol{x}_{i}^{t}):\boldsymbol{x}_{i}^{t}\in\boldsymbol{X}_{i}\})|<\varepsilon.$ (3)
Theorem 1 defines the generic pipeline of a MIL, which consists of three main parts: (i) The function $f$ is a feature extractor that projects the input instances into $L$-dimensional vector embeddings $\tilde{\boldsymbol{X}}_{i}$. (ii) $\sigma$ is known as the MIL pooling function that aggregates instance vector embeddings into a single vector. It should be noted that the original MIL pooling function $\sigma$ should be permutation-invariant (Ilse et al., 2018). (iii) $g$ denotes the bag-level classifier (e.g., a linear classifier) that maps the vector embedding after applying MIL pooling to a bag-level probability prediction $\hat{y}_{i}\in[0,1]$.
Mutli-Class MTSC. A multi-class time series classification with a total of $C$ classes can be performed as several one-vs-rest binary MIL without violating its assumption:
$y_{i,c}=\left\{\begin{array}[]{l}0,\text{ iff }\sum_{t=1}^{T}y_{i,c}^{t}=0,\ y^{t}_{i,c}\in\{0,1\}\\ 1,\text{ otherwise,}\end{array}\right.$ (6)
where $y_{i,c}^{t}=1$ denotes a time point with significant contribution to class $c\in\{1,\cdots,C\}$. The final bag-level prediction $y_{i}$ for a bag is computed as the class with the highest probability:
$\hat{y}_{i}=\operatorname{argmax}\limits_{c}\ \hat{y}_{i,c},$ (7)
which is consistent with the one-vs-rest scheme.
Remark 1.
The MIL in Theorem 1 fails to model the temporal ordering among time points within a time series.
Remark 1 arises from the symmetric (permutation-invariant) property of the MIL pooling function $\sigma$. The function $\sigma$ remains the same for every permutation of the instances within a bag, thereby neglecting the temporal ordering between time points (instances) in time series modeling. This hinders the direct translation of classic MIL into MTSC tasks. To address this limitation, we propose a time-aware MIL pooling in Section 3.3.
3.3 Time-Aware MIL Pooling for MTSC
From an entropy perspective, understanding permutation-variant properties in time series can be quite insightful. As discussed in Section 3.2, we assume that each time point (instance) $\boldsymbol{x}_{i}^{j}$ in a time series (bag) is a realization of a random variable $\Theta^{j}$ conditioned by a time index $t^{j}$. The resulting bag can be represented as a random variable $\boldsymbol{X}=\{(\Theta^{1}|t^{1}),\cdots,(\Theta^{T}|t^{T})\}$, where $t^{j}\neq t^{k},\forall j\neq k$. Likewise, $\boldsymbol{Y}$ denotes a random variable for bag label. Here, we use notation $\Theta^{j}$ for an instance to distinguish a random variable (often as uppercase) and its realization (often as lowercase). It is fair to construct a general assumption: $p(\Theta^{j}|t^{j})\neq p(\Theta^{j}|t^{k})$, which indicates an instance varies when it is presented in different locations.
Proposition 2.
Shuffling the time points within a time series potentially disrupts its predictability. This means, under the general assumption, the entropy before and after shuffling typically differs, i.e., the equality:
$H(\cdots,(\Theta^{j}|t^{j}),\cdots)=H(\cdots,(\Theta^{j}|t^{\bar{j}}),\cdots),$
does not always hold. Here, $t^{\bar{1}},\cdots,t^{\bar{T}}$ are sampled from the set $\{t^{1},\cdots,t^{T}\}$ without replacement. The right term denotes the time series after randomly shuffling.
Figure 3: The block entropy in Shakespeare’s Sonnets with varying shuffling rates, where the higher shuffling rates result in higher block entropy.
Please refer to Appendix C for its proof. Proposition 2 implies that random permutation of time points within a regular time series potentially increases its uncertainty in terms of entropy. Since it is challenging to directly compute the entropy for high-dimensional continuous-valued variables from the observation, we illustrate this fact by using a supportive example in text sequence (Shakespeare’s Sonnets111 https://shakespeares-sonnets.com/Archive/allsonn.htm.). The feasible domain of the text sequence is a discrete 1D domain, and its uncertainty can be measured directly by the Shannon block entropy (Shannon, 1951). Please refer to Appendix D for more detail about Shannon block entropy. As shown in Fig. 3, the block entropy increases as the shuffling rate increases. We will demonstrate the importance of not just modeling the random permutation but also accounting for the temporal correlation between instances.
Theorem 3.
Modeling the temporal correlation between instances lowers the complexity of developing a good classifier, which is presented by class-conditioned entropy:
$H(\boldsymbol{X}_{c}|\boldsymbol{Y})\leq H(\boldsymbol{X}_{nc}|\boldsymbol{Y}),$ (8)
where $H(\boldsymbol{X}_{c}|\boldsymbol{Y})=H((\Theta^{1}|t^{1}),\cdots,(\Theta^{T}|t^{T})|\boldsymbol{Y})$ is derived under modeling the correlation among instances within a bag while $H(\boldsymbol{X}_{nc}|\boldsymbol{Y})=\sum_{i}H((\Theta^{i}|t^{i})|\boldsymbol{Y})$ is derived under the assumption that instances are independent and identically distributed.
Proof.
For convenience, we denote $(\Theta^{j}|t^{j})$ by $\Lambda^{j}$.
$\begin{split}H(\Lambda^{1},\cdots,\Lambda^{T}|\boldsymbol{Y})&=H(\Lambda^{T}|\Lambda^{T-1},\cdots,\Lambda^{1},\boldsymbol{Y})\\ &+H(\Lambda^{T-1}|\Lambda^{T-2},\cdots,\Lambda^{1},\boldsymbol{Y})\\ &+\cdots\\ &+H(\Lambda^{1}|\boldsymbol{Y})\end{split}$ (9)
Since $H(\Lambda^{i}|\Lambda^{i-1},\cdots,\Lambda^{1},\boldsymbol{Y})\leq H(\Lambda^{i}|\boldsymbol{Y})$, which indicates knowing more information lowers the uncertainty:
$H(\Lambda^{1},\cdots,\Lambda^{T}|\boldsymbol{Y})\leq\sum_{i}H(\Lambda^{i}|\boldsymbol{Y})$ (10)
Remark 2.
The conditional entropy $H(\boldsymbol{X}|\boldsymbol{Y})$ measures the uncertainty of the bag feature $\boldsymbol{X}$ given that the bag-level class label $\boldsymbol{Y}$ is known. In the context of classification, it quantifies the spread of features within each class. A high value of $H(\boldsymbol{X}|\boldsymbol{Y})$ indicates that the features belonging to the same class can vary significantly. This suggests that the features are not clustered tightly but are spread out. In contrast, a lower $H(\boldsymbol{X}|\boldsymbol{Y})$ suggests that the features from the same class are more homogeneous, exhibiting less variability. This homogeneity can make it easier to classify instances since the features within each class are more consistent, potentially resulting in a simpler decision boundary.
Theorem 3 immediately implies the benefit of modeling temporal permutation and correlation between instances and provides a generic formulation of time-aware MIL pooling. The realization of Theorem 3 can be achieved by a transformer with the unique token mechanism. First, transformers help tackle the conditional entropy $H(\Lambda^{t}|\Lambda^{t-1},\cdots,\Lambda^{1},\boldsymbol{Y})$ in Theorem 3 by employing a class token $\boldsymbol{x}_{i}^{\operatorname{cls}}$. The yielded tokenized bag of instances is $\boldsymbol{X}_{i}^{\operatorname{cls}}=\{\boldsymbol{x}_{i}^{\operatorname{cls}},\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\}$. Second, we propose a novel positional encoding in our transformer-based pooling through the lens of wavelet theory to further capture the multi-scale time-frequency ordering relationship among instances.
Temporal Correlation as Self-Attention. The self-attention mechanism (Vaswani et al., 2017b) is proposed to capture mutual information between time points. In the context of MIL, we use multi-head self-attention (MHSA) to model the sequential correlation between instances:
$\operatorname{MHSA}(\boldsymbol{X}_{i}^{\operatorname{cls}})=[\operatorname{head}_{1},\cdots,\operatorname{head}_{H}]\boldsymbol{W}_{0}\vspace{-0.6cm}$ (11)
where:
$\operatorname{head}_{h}=\operatorname{Attention}(\boldsymbol{X}_{i}^{\operatorname{cls}}\boldsymbol{W}_{h}^{Q},\boldsymbol{X}_{i}^{\operatorname{cls}}\boldsymbol{W}_{h}^{K},\boldsymbol{X}_{i}^{\operatorname{cls}}\boldsymbol{W}_{h}^{V})$
$\operatorname{Attention}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\operatorname{softmax}(\boldsymbol{Q}\boldsymbol{K}^{\mathbf{T}}/\sqrt{d_{k}})\boldsymbol{V},$ (12)
where $\boldsymbol{W}_{0}$, $\boldsymbol{W}_{h}^{Q}$, $\boldsymbol{W}_{h}^{K}$, $\boldsymbol{W}_{h}^{V}$ are trainable parameters. It is noteworthy that after the transformer blocks, we only pass the class token to the bag-level classifier $g$ to make a prediction. However, standard self-attention has quadratic time and memory complexity $\mathcal{O}(T^{2})$ w.r.t. the number of instances in a bag ($T$). Recent advances (Xiong et al., 2021; Wang et al., 2020; Shen et al., 2021) in self-attention studies have reduced the quadratic complexity to approximately linear. Specifically, we use the approximation of self-attention proposed by (Xiong et al., 2021) in our implementation to reduce the complexity of the proposed TimeMIL.
Figure 4: The proposed learnable wavelet positional encoding: First, wavelet transform is performed for the input signal (by excluding the class token) with each wavelet basis (Eq. 17). Second, the signals are aggregated in the wavelet domain by a summation (Eq. 13). In the case of $n_{w}=3$, we use 3 learnable wavelet bases ($\boldsymbol{\Psi}_{1},\boldsymbol{\Psi}_{2},\boldsymbol{\Psi}_{3}$) to model changing frequency and time scales.
Wavelet Positional Encoding. The classic transformers use Sinusoidal positional encoding to capture the relative ordering in a time series, as self-attention does not take the temporal ordering of time points into account. The disadvantage of Sinusoidal positional encoding is that it is pre-defined and non-learnable. Importantly, the Sinusoidal positional encoding is independently generated away from the input context; hence, it cannot capture both the time and frequency information of time series. To address this limitation, we propose a learnable wavelet positional encoding in a conditional positional encoding fashion (Chen et al., 2023; Chu et al., 2023):
$\vspace{-0.2cm}\operatorname{WPE}(\boldsymbol{X}_{i})=\sum_{j=1}^{n_{\operatorname{W}}}\Phi(\boldsymbol{X}_{i},\boldsymbol{\Psi}_{j}),\vspace{-0.1cm}$ (13)
where $n_{\operatorname{W}}$ denotes the number of wavelet basis, which is empirically set to 3 in this paper. $\{\boldsymbol{\Psi}_{1},\cdots,\boldsymbol{\Psi}_{n_{\operatorname{W}}}\}$ are learnable wavelet kernels with $\boldsymbol{\Psi}_{j}=\{\psi_{a_{j1},b_{j1}}(t),\cdots,\psi_{a_{jL},b_{jL}}(t)\}$ where $\psi$ is mother wavelet, which is chosen to be the Mexican hat in our experiments. The Gabor-Heisenberg limit (Gabor, 1946), which is the uncertainty principle in the time-frequency version, states that it is impossible to precisely determine both the time and frequency of a signal, simultaneously. In the context of the wavelet transform, this principle implies a trade-off: the higher the resolution needed in time, the lower the resolution becomes in frequency. This implies that the careful selection of time-frequency resolution is crucial for effectively characterizing different signals. Hence, we learn the scaling and translation parameters $\{(a_{j1},b_{j1}),\cdots,(a_{jL},b_{jL})\}$ to form the wavelet basis from mother wavelet. $\Phi(\cdot)$ is the channel-wise wavelet transform, which can be formulated as convolving the input signal with the wavelet kernels:
$\Phi(\tilde{\boldsymbol{X}}_{i},\boldsymbol{\Psi}_{j})=\left[\begin{array}[]{c}\tilde{\boldsymbol{X}}_{i1}\circledast\psi_{a_{j1},b_{j1}}\\ \vdots\\ \tilde{\boldsymbol{X}}_{iL}\circledast\psi_{a_{jL},b_{jL}}\end{array}\right]^{\top}\in\mathbb{R}^{n\times L}.$ (17)
3.4 Interpretability
The proposed time-aware MIL is naturally interpretable due to its ability to localize time points of interest within a time series. This is achieved by hacking into the attention map of the transformer layers in the proposed time-aware MIL pooling. The attention map (refer to Eq. 3.3) measures the importance of each instance $\boldsymbol{x}_{i}^{t}$ in the series $\boldsymbol{X}_{i}$ in the MIL pooling:
$\boldsymbol{A}_{i}=\operatorname{softmax}\left(\frac{\boldsymbol{x}_{i}^{\operatorname{cls}}\boldsymbol{W}_{h}^{Q}\ \boldsymbol{X}_{i}\boldsymbol{W}_{h}^{K}}{\sqrt{d_{k}}}\right),$ (18)
where $\operatorname{softmax}$ is performed over the time dimension. $\boldsymbol{A}_{i}\in\mathbb{R}^{T}$ is the importance weights of all instances, with its $t$-th element corresponding to the importance of the time point $\boldsymbol{x}_{i}^{t}$. We only use the class token to calculate the importance weight, as it determines the most relevant time points that characterize a certain class label in MTSC tasks.
Algorithm 1 Time-aware MIL (Forward Propagation)
0:  Input sequence $\boldsymbol{X}_{i}=\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}$. A neural network contains: $f$: feature extractor, $g$: bag-level classifier, $WPE$ bases, and $\boldsymbol{x}^{\operatorname{cls}}$: class token. Two Transfomer blocks.
0:  Predicted label $\hat{y}_{i}$.
1:  Get embedding: $\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}\leftarrow f(\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\})$MIL pooling from here (Section 3.3)
2:  Init class token for the bag: $\boldsymbol{x}_{i}^{\operatorname{cls}}\leftarrow\boldsymbol{x}^{\operatorname{cls}}$.
3:  Add class token: $\boldsymbol{X}_{i}^{\operatorname{cls}}=\boldsymbol{x}_{i}^{\operatorname{cls}}\cup\boldsymbol{X}_{i}$.
4:  for j=1:2 do
5:     PE: $\boldsymbol{X}_{i}\leftarrow\boldsymbol{X}_{i}+WPE_{j}(\boldsymbol{X}_{i})$. # $WPE_{j}(\boldsymbol{X}_{i})$ is from Eq. 10. Here, $\boldsymbol{X}_{i}$ is the part of $\boldsymbol{X}_{i}^{\operatorname{cls}}$.
6:     $\boldsymbol{X}_{i}^{\operatorname{cls}}\leftarrow Transformer_{j}(\boldsymbol{X}_{i}^{\operatorname{cls}})$.
7:  end for#Bag-level classification from here
8:  $\hat{y}_{i}\leftarrow g(\boldsymbol{x}_{i}^{\operatorname{cls}})$. #$\boldsymbol{x}_{i}^{\operatorname{cls}}$ is obtained from $\boldsymbol{X}_{i}^{\operatorname{cls}}$.
Table 1: Results of Group 1 Experiments. Comparison with the recent state-of-the-art MTSC methods on 26 datasets. The best results are highlighted by bold and the second best are highlighted by underline. ’N/A’ in the table denotes the corresponding method was unable to obtain results due to memory or computational limitations (Liu et al., 2023).
Datasets/Methods ED-1NN DTW-1NN-I DTW-1NN-D
MLSTM-FCN
Neur. Net.’19
ShapeNet
AAAI’21
WEASEL+MUSE
arxiv’2017
TapNet
AAAI’20
OS-CNN
ICLR’22
MOS-CNN
ICLR’22
TodyNet
arxiv’23
Ours
ArticularyWordRecognition 0.970 0.980 0.987 0.973 0.987 0.990 0.987 0.988 0.991 0.987 0.990
AtrialFibrillation 0.267 0.267 0.200 0.267 0.400 0.333 0.333 0.233 0.183 0.467 0.733
BasicMotions 0.675 1.000 0.975 0.950 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Cricket 0.944 0.986 1.000 0.917 0.986 1.000 0.958 0.993 0.990 1.000 1.000
DuckDuckGeese 0.275 0.550 0.600 0.675 0.725 0.575 0.575 0.540 0.615 0.580 0.780
EigenWorms 0.550 0.603 0.618 0.504 0.878 0.890 0.489 0.414 0.508 0.840 0.823
Epilepsy 0.667 0.978 0.964 0.761 0.987 1.000 0.971 0.980 0.996 0.971 1.000
EthanolConcentration 0.293 0.304 0.323 0.373 0.312 0.133 0.323 0.240 0.415 0.350 0.407
ERing 0.133 0.133 0.133 0.133 0.133 0.430 0.133 0.881 0.915 0.915 0.956
FaceDetection 0.519 0.513 0.529 0.545 0.602 0.545 0.556 0.575 0.597 0.627 0.698
FingerMovements 0.550 0.520 0.530 0.580 0.589 0.490 0.530 0.568 0.568 0.570 0.670
HandMovementDirection 0.279 0.306 0.231 0.365 0.338 0.365 0.378 0.443 0.361 0.649 0.487
Handwriting 0.371 0.509 0.607 0.286 0.451 0.605 0.357 0.668 0.677 0.436 0.482
Heartbeat 0.620 0.659 0.717 0.663 0.756 0.727 0.751 0.489 0.604 0.756 0.815
Libras 0.833 0.894 0.872 0.856 0.856 0.878 0.850 0.950 0.965 0.850 0.972
LSST 0.456 0.575 0.551 0.373 0.590 0.590 0.568 0.413 0.521 0.615 0.690
MotorImagery 0.510 0.390 0.500 0.510 0.610 0.500 0.590 0.535 0.515 0.640 0.720
NATOPS 0.860 0.850 0.883 0.889 0.883 0.870 0.939 0.968 0.951 0.972 0.994
PenDigits 0.973 0.939 0.977 0.978 0.977 0.948 0.980 0.985 0.983 0.987 0.600
PEMS-SF 0.705 0.734 0.711 0.699 0.751 N/A 0.751 0.760 0.764 0.780 0.931
PhonemeSpectra 0.104 0.151 0.151 0.110 0.298 0.190 0.175 0.299 0.295 0.309 0.311
RacketSports 0.868 0.842 0.803 0.803 0.882 0.934 0.868 0.877 0.929 0.803 0.908
SelfRegulationSCP1 0.771 0.765 0.775 0.874 0.782 0.710 0.652 0.835 0.829 0.898 0.898
SelfRegulationSCP2 0.483 0.533 0.539 0.472 0.578 0.460 0.550 0.532 0.510 0.550 0.639
StandWalkJump 0.200 0.333 0.200 0.067 0.533 0.333 0.400 0.383 0.383 0.467 0.733
UWaveGestureLibrary 0.881 0.869 0.903 0.891 0.906 0.916 0.894 0.927 0.926 0.850 0.900
Ours 1-to-1-Wins 25 23 22 25 22 16 24 22 19 20 -
Ours 1-to-1-Draws 0 1 1 0 1 4 1 1 1 3 -
Ours 1-to-1-Losses 1 2 3 1 3 5 1 3 6 3 -
Average accuracy ($\uparrow$) 0.568 0.622 0.626 0.597 0.684 0.656 0.637 0.672 0.692 0.726 0.774
Total best accuracy ($\uparrow$) 0 1 1 0 1 5 1 2 4 5 18
Average Rank ($\downarrow$) 9.154 7.904 7.231 7.865 4.731 5.900 6.500 5.442 4.692 4.058 2.327
Table 2: Results of Group 2 Experiments. Comparison with the recent state-of-the-art general time analysis frameworks on 10 datasets.
Datasets/Methods
LSTNet
SIGIR’18
LSSL
ICLR’22
Rocket
DMKD’18
SCINet
NeurIPS’22
MICN
ICLR’23
TimesNet
ICLR’23
Dlinear
AAAI’23
LightTS
arxiv’22
MTS-Mixer
arxiv’23
Rlinear
arxiv’23
RMLP
arxiv’23
FEDformer
ICML’22
Flowformer
ICML’22
Crossformer
ICLR’23
PatchTST
ICLR’23
ModernTCN
ICLR’24
Ours
EthanolConcentration 0.399 0.311 0.452 0.344 0.353 0.357 0.362 0.297 0.338 0.289 0.313 0.312 0.338 0.380 0.328 0.363 0.407
FaceDetection 0.657 0.667 0.647 0.689 0.652 0.686 0.680 0.675 0.702 0.656 0.673 0.660 0.676 0.687 0.683 0.708 0.698
Handwriting 0.258 0.246 0.588 0.236 0.255 0.321 0.270 0.261 0.260 0.281 0.300 0.280 0.338 0.288 0.296 0.306 0.487
Heartbeat 0.771 0.727 0.756 0.775 0.747 0.780 0.751 0.751 0.771 0.726 0.727 0.737 0.776 0.776 0.749 0.772 0.815
JapaneseVowels 0.981 0.984 0.962 0.960 0.946 0.984 0.962 0.962 0.943 0.959 0.959 0.984 0.989 0.991 0.975 0.988 0.995
PEMS-SF 0.867 0.861 0.751 0.838 0.855 0.896 0.751 0.884 0.809 0.827 0.839 0.809 0.860 0.859 0.893 0.891 0.931
SelfRegulationSCP1 0.840 0.908 0.908 0.925 0.860 0.918 0.873 0.898 0.917 0.911 0.921 0.887 0.925 0.921 0.907 0.934 0.898
SelfRegulationSCP2 0.528 0.522 0.533 0.572 0.536 0.572 0.505 0.511 0.550 0.561 0.510 0.544 0.561 0.583 0.578 0.603 0.639
SpokenArabicDigits 1.000 1.000 0.712 0.981 0.971 0.990 0.814 1.000 0.974 0.965 0.976 1.000 0.988 0.979 0.983 0.987 1.000
UWaveGestureLibrary 0.878 0.859 0.944 0.851 0.828 0.853 0.821 0.803 0.823 0.825 0.838 0.853 0.866 0.853 0.858 0.867 0.900
Ours 1-to-1-Wins 9 8 6 9 10 9 10 8 8 9 9 9 9 9 9 8 -
Ours 1-to-1-Draws 1 1 0 0 0 0 0 2 0 0 0 1 0 0 0 0 -
Ours 1-to-1-Losses 0 1 4 1 0 1 0 0 2 1 1 0 1 1 1 2 -
Average accuracy ($\uparrow$) 0.718 0.709 0.725 0.717 0.700 0.736 0.679 0.704 0.709 0.700 0.706 0.707 0.732 0.732 0.725 0.742 0.777
Total best accuracy ($\uparrow$) 1 1 3 0 0 0 0 1 0 0 0 1 0 0 0 2 5
Average Rank ($\downarrow$) 8.85 10.4 9.5 8.9 13.2 5.35 12.7 11.2 10.85 13 11.55 10.75 5.9 5.9 8.1 4 2.85
4 Experiments
4.1 Setup and Baselines
We use the UEA benchmark datasets to validate the superiority of the proposed TimeMIL. These datasets have various lengths, dimensions, and training/test splits. Please refer to Appendix A for the details of these datasets. We conduct two groups of experiments as follows.
Group 1 experiments. Following (Liu et al., 2023; Li et al., 2021b), this group of experiments is conducted on selected 26 equal-length datasets from UEA to compare with the recent strong baseline methods of multivariate time series classification, including: TodyNet (Liu et al., 2023), OS-CNN and MOS-CNN (Tang et al., 2022), ShapeNet (Li et al., 2021b), TapNet (Zhang et al., 2020), WEASEL+MUSE (Schäfer & Leser, 2017), WLSTM-FCN (Karim et al., 2019). We also include several well-known traditional methods based on distance and the nearest neighbor classifier, including ED-1NN, DTW-1NN-I, and DTW-1NN-D. Please refer to (Liu et al., 2023) or Appendix F for the details of these three baselines.
Group 2 experiments. Following (donghao & wang xue, 2024; Wu et al., 2023), this group of experiments is conducted on selected 10 UEA datasets to compare with the recent strong methods in the general time series analysis, including: ModernTCN (donghao & wang xue, 2024), PatchTST (Nie et al., 2023), Crossformer (Zhang & Yan, 2023), Flowformer (Wu et al., 2022), FEDformer (Zhou et al., 2022), Rlinear and RMLP (Li et al., 2023b), MTS-Mixer (Li et al., 2023c), LightTS (Zhang et al., 2022b), Dlinear (Zeng et al., 2023), TimesNet (Wu et al., 2023), MICN (Wang et al., 2023), SCINet (Liu et al., 2022), Rocket (Dempster et al., 2020), LSSL (Gu et al., 2022), and LSTNET (Lai et al., 2018a).
Implementation details. The previous benchmark results of all baseline methods are taken from their papers, and the same training setting is used. In the implementation of our proposed model, we adopt the model in (Ismail Fawaz et al., 2020) as our backbone, where the output dimension $L$ is fixed to 128. Refer to Appendix F for details.
Evaluation metrics. Following (Liu et al., 2023; Li et al., 2021b), we evaluate the performance of our proposed method and other methods by computing the accuracy, average accuracy, average rank, and the number of pair-wise Wins/Draws/Losses.
4.2 Main Experimental Results
The proposed method demonstrates superiority over other recently proposed competing methods in both Group 1 and Group 2 experiments (refer to Table 1 and 2).
Group 1 results. We obtain a 77.4% average accuracy on all 26 MTSC datasets, surpassing other methods on 18 datasets and achieving an average rank of 2.327 out of a total of 11 methods. Specifically, the proposed method outperforms the second-best methods on each dataset by an average of 4.8% in accuracy and reduces the performance rank by 1.73. This performance gain is even more substantial in those challenging datasets. Specifically, compared to the second-best performing methods, our method improves the average accuracy by 26%, 20%, 15%, 8%, 8%, 7%, and 7% on AtrialFibrillation, StandWalkJump, PEMS-SF, FingerMovements, MotorImagery, FaceDetection, and LSST datasets, respectively.
Group 2 results. The proposed method also achieved superior performance on the Group 2 datasets compared to other methods. Specifically, our method achieved a 77.7% average accuracy, surpassing the other methods in 5 out of 10 datasets and achieving an average rank of 2.85 out of a total of 17 methods. Notably, the proposed method outperforms the recent state-of-the-art ModernTCN (donghao & wang xue, 2024) by 3.5% in average accuracy and 1.15 in average rank.
Table 3: Comparison of different MIL pooling (Left) and positional encoding (Right) with Group 1 experiments.
MIL Pooling Accuracy
Mean 0.715
Max 0.719
Attention 0.739
Conjunctive 0.746
Time-aware (Ours) 0.774
Positional Encoding Accuracy
ABMIL TimeMIL
None 0.739 0.761
Sinusoidal 0.741 0.763
WPE (Ours) 0.745 0.774
Figure 5: Exemplary attention maps learned in TimeMIL using different datasets (rows) including synthetic dataset, StandWalkJump dataset, and AtrialFibrillation dataset, featuring distinct patterns of interest (columns). TimeMIL accurately localized patterns of interest.
Figure 6: Decision boundary learned in fully supervised method (Left) versus TimeMIL (Right) using the synthetic dataset.
4.3 Ablation on Model Design Variants
Effectiveness of TimeMIL pooling. We compare the performance of the proposed TimeMIL with other commonly used MIL pooling methods, including MeanMIL, MaxMIL, ABMIL (Ilse et al., 2018) and the most recent ConjunctiveMIL (Early et al., 2024). We observe that learnable pooling methods (i.e., TimeMIL, ABMIL, and ConjunctiveMIL) show superior performance over non-parametric MIL pooling methods (i.e., MeanMIL and MaxMIL) (Table 3 Left). Notably, the proposed TimeMIL outperforms ABMIL and ConjunctiveMIL by 3% and 2.8% in terms of average accuracy, respectively. This supports our initial claim that modeling dependencies between time points is beneficial for MTSC.
Effectiveness of wavelet positional encoding. We compare the proposed WPE with commonly used Sinusoidal PE (Vaswani et al., 2017a). We observe that adding PE generally improves the performance of both ABMIL and TimeMIL (see Table 3 Right), which aligns with our hypothesis that incorporating PE into MIL can better model the ordering within time points and hence lower the classification error. However, we do not observe a significant performance gain ($\sim$ 0.27%) by adding a sinusoidal PE for both ABMIL and TimeMIL. On the contrary, the addition of the proposed WPE improves ABMIL and TimeMIL by 0.6% and 1%, respectively. This may be attributed to the fact that it is challenging for predefined sinusoidal PE to capture the changing of frequencies over time within a time series. While the proposed WPE better models these time-frequency changes.
4.4 The effectiveness of weakly supervised learning
To validate the effectiveness of the proposed weakly supervised learning scheme, we provide an in-depth comparison between the proposed TimeMIL and traditional fully supervised methods. The results are shown in Fig. 5 and 6.
Decision boundary. We visualize the decision boundary learned in weakly supervised TimeMIL and those learned in the fully supervised method using the synthetic dataset. For the implementation of the supervised method, we maintain the same feature extractor but replace the TimeMIL pooling with a supervised classifier (Ismail Fawaz et al., 2020). For visualization of the decision boundary, we projected the time series data onto 2D space by performing PCA on the latent space (i.e., the feature embeddings after applying the feature extractor). More details of the synthetic dataset and visualizing the decision boundary can be found in Appendix H.
We observe that although the fully supervised method could differentiate positive and negative time series, there is not an apparent margin in the decision boundary between the positive and negative time series (see Fig. 6 (Left)). This is because the negative instances in positive bags (time series) resemble those negative instances in negative bags. Meanwhile, the positive instances in positive bags are typically less than negative instances. In contrast, the decision boundary of the weakly supervised TimeMIL shows a distinct decision boundary with large margins between the positive and negative instances (Fig. 6 (Right)). This unique feature of the proposed TimeMIL provides an instance differentiation between positive and negative time series, offering a precise localization of positive instances.
Inherent interpretability. The instance-level decision-making mechanism makes the proposed TimeMIL inherently interpretable. Leveraging this, we can obtain the importance score (Eq. 18). TimeMIL accurately localized patterns of interest (i.e., positive instances) in the UEA dataset (Fig. 5; $2^{nd}$ and $3^{rd}$ row) and synthetic dataset (Fig. 5; $1^{st}$ row). This supports our initial hypothesis that time points of interest in positive time series are typically sparse and localized, making the weakly supervised TimeMIL a natural choice for MTSC.
5 Conclusion
In this work, we introduce TimeMIL, a weakly supervised MIL framework designed for multivariate time series classification. The proposed method, from the perspective of weakly supervised learning, offers a better capability to characterize the decision boundary for MTSC than commonly used fully supervised methods. As a result, TimeMIL demonstrates superiority over 26 methods in 28 datasets and illustrates impressive interpretability by accurately localizing patterns of interest.
6 Limitation
Currently, our work primarily focuses on modeling the sparsity and locality of time points, and we have not yet extended the interpretability pooling to the channel level. Our work may also be limited by the case that a time point fuses multiple variates that represent potential delayed events or measurements. We envision that a major difference in expanding to the channel level would involve designing a channel-based positional encoding. This will be a topic of exploration in our future work.
7 Impact Statements
This paper presents a novel TimeMIL framework to model the temporal correlation and ordering, leveraging a tokenized transformer with a specialized learnable wavelet positional token. The potential societal consequences of TimeMIL can be summarised three-fold: (i) Theoretical View: Commencing with a viewpoint in weakly supervised learning, we are the first to formally introduce MIL into time series-related tasks. We rigorously examine feasibility from an information-theoretic perspective. In response to the various limitations of MIL in MTSC, we propose corresponding modules to preserve the intrinsic temporal correlation and ordering properties within time series. We draw connections between the MIL and MTSC. This provides an improved justification, relating it to existing MTSC methods. (ii) Flexibility: Our framework significantly expands the MTSC applications. For example, future work involving financial forecasting, predictive maintenance, anomaly detection, healthcare monitoring, financial forecasting, environmental monitoring, and speech and signal processing, among others, can all be incorporated within this framework. (iii) Interpretability: Compared with the previously existing methods, TimeMIL could provide the interpretable interest of the pattern of the network. By visualizing the points of time series that strongly influence predictions, TimeMIL can assist in ensuring model robustness. This can help in identifying vulnerabilities and reducing the risk of adversarial attacks. Meanwhile, Interpretability contributes to the explainability of MTSC, making it easier for researchers, practitioners, and stakeholders to comprehend/validate how and why a model makes specific predictions.
Appendix A UEA Datasets Detail
The detail of all 30 datasets is provided in Table 4. It should be noteworthy that the datasets JapaneseVowels and SpokenArabicDigits used in the Group 2 Experiment originally have varied lengths of sequences. We pre-process it following (Wu et al., 2023), where we pad them to 29 and 93, respectively.
Table 4: Dataset Summary
Dataset Train Size Test Size Dimensions Length Classes
ArticularyWordRecognition 275 300 9 144 25
AtrialFibrillation 15 15 2 640 3
BasicMotions 40 40 6 100 4
CharacterTrajectories 1422 1436 3 182 20
Cricket 108 72 6 1197 12
DuckDuckGeese 60 40 1345 270 5
EigenWorms 128 131 6 17984 5
Epilepsy 137 138 3 206 4
EthanolConcentration 261 263 3 1751 4
ERing 30 30 4 65 6
FaceDetection 5890 3524 144 62 2
FingerMovements 316 100 28 50 2
HandMovementDirection 320 147 10 400 4
Handwriting 150 850 3 152 26
Heartbeat 204 205 61 405 2
JapaneseVowels 270 370 12 29 (max) 9
Libras 180 180 2 45 15
LSST 2459 2466 6 36 14
InsectWingbeat 30000 20000 200 78 10
MotorImagery 278 100 64 3000 2
NATOPS 180 180 24 51 6
PenDigits 7494 3498 2 8 10
PEMS-SF 267 173 963 144 7
Phoneme 3315 3353 11 217 39
RacketSports 151 152 6 30 4
SelfRegulationSCP1 268 293 6 896 2
SelfRegulationSCP2 200 180 7 1152 2
SpokenArabicDigits 6599 2199 13 93 (max) 10
StandWalkJump 12 15 4 2500 3
UWaveGestureLibrary 120 320 3 315 8
Appendix B More detail of Theorem 1
We provide an intuition behind this from the perspective of Hausdorff distance, which is defined to measure the difference between two sets from the same feasible domain. Let $(\mathcal{X},d)$ be a metric space, for each pair of non-empty sets ($\boldsymbol{X},\boldsymbol{X}^{\prime})\subset\mathcal{X}$, their Hausdorff distance $d_{\mathrm{H}}$ is computed as
[…]
$\sup_{\boldsymbol{x}_{j}^{t}\in\boldsymbol{X}_{j}}\inf_{\boldsymbol{x}_{i}^{t^{\prime}}\in\boldsymbol{X}_{i}}d(\boldsymbol{x}_{j}^{t},\boldsymbol{x}_{i}^{t^{\prime}})\},$ (19)
where $\sup$ and $\inf$ denote the supremum operator and the infimum operator, respectively. Eq. B implies that Hausdorff distance measures the furthest distance of traveling from a certain point in a set to its nearest point in the other set under the worst-case scenario. Hence, Hausdorff distance can only measure the distance of time points across different sets and fails to model the time ordering information.
Appendix C Proof of Proposition 2
Proof.
The existence of equality can be easily proved by assuming the time series is always a constant value, which, regardless of permutation, the entropy is 0. The existence of inequality can be proved by contradiction. Suppose a sequence with a length of 2, presenting two tests $\{(\Theta^{1}|t^{1}),(\Theta^{2}|t^{2})\}$ for a product. Suppose the probability of passing each test obeys $(\Theta^{i})\sim Bernoulli(p=e^{-3+i})$. Suppose the ordering information says the test with index $t^{2}$ occurs if the product passes the test with index $t^{1}$. we have $p(0,0)=1-e^{-2},p(0,1)=0,p(1,0)=e^{-2}(1-e^{-1}),p(1,1)=e^{-1}e^{-2}$.
$H(\{(\Theta^{1}|t^{1}),(\Theta^{2}|t^{2})\})$ (20)
$=$ $-(1-e^{-2})\log(1-e^{-2})-0\log 0-e^{-2}(1-e^{-1})\log(e^{-2}(1-e^{-1}))-e^{-1}e^{-2}\log(e^{-1}e^{-2})$
[…]
After permuting sequence, $p(0,0)=1-e^{-1},p(0,1)=0,p(1,0)=e^{-1}(1-e^{-2}),p(1,1)=e^{-1}e^{-2}$. The entropy of the permuted sequence is presented as,
$H(\{(\Theta^{1}|t^{2}),(\Theta^{2}|t^{1})\})$ (21)
$=$ $-(1-e^{-1})\log(1-e^{-1})-0\log 0-e^{-1}(1-e^{-2})\log(e^{-1}(1-e^{-2}))-e^{-1}e^{-2}\log(e^{-1}e^{-2})$
[…]
which are apparently different, and the existence of inequality is proved.
Appendix D Block Entropy
Block entropy, also known as N-gram entropy, is used to measure the uncertainty of a sequence by (Shannon, 1951). Suppose a list of overlapping blocks is generated via a sliding window with a size of $n$, where the $j$-th block is $B_{j}^{(n)}=\left(X_{j},\ldots,X_{j+n-1}\right)$. Suppose the set of all appearance of blocks denotes $\left\{b_{1}^{(n)},\cdots,b_{i}^{(n),\cdots}\right\}$. For example, in a sequence $AAABBCD$, the set of all possible blocks is $\{AA,AB,BB,BC,CD\}$. Then, the block entropy is defined by
$H_{n}=-\sum_{i=1}^{L^{n}}p\left(b_{i}^{(n)}\right)\log\left(p\left(b_{i}^{(n)}\right)\right),$ (22)
where $p\left(b_{i}^{(n)}\right)$ denotes the probability of appearance of the block sequence $b_{i}^{(n)}$. We set $n=2$ in our experiment.
Appendix E Background of Wavelet Transform
Mathematically, a wavelet basis $\psi_{a,b}$ is a generated by scaling $a$ and translations $b$ of a single function named mother wavelet $\psi\in L^{2}(\mathbb{R})$, where $L^{2}(\mathbb{R})$ denotes the Hilbert space of square integrable functions,
$\psi_{a,b}(t)=\frac{1}{\sqrt{|a|}}\psi\left(\frac{x-b}{a}\right).$ (23)
It is noteworthy that the basis $\psi_{a,b}$ is a Hilbert basis, which implies every basis is orthogonal with each other,
[…]
This also ensures that different wavelet basis are exploring diverse context of the signal. The Continuous Wavelet Transform (CWT) of a 1D signal $f(t)$ is defined by
[…]
Theorem 4.
(Gabor, 1946) Uncertainty principles in the time-frequency version (also known as the Gabor-Heisenberg limit ):
$\sigma_{t}\cdot\sigma_{f}\geq\frac{1}{4\pi},$ (26)
where $\sigma_{t}$ and $\sigma_{f}$ denote the measured time and frequency standard deviations, respectively.
We show more diverse learned Wavelet kernel in Fig. 7.
Figure 7: The learned wavelet kernel from MTSC tasks.
Appendix F Implementation Detail
F.1 Baselines
ED-1NN, DTW-1NN-I, and DTW-1NN-D are most popular baselines for MTSC: (i) ED-1NN: It applies the nearest neighbor classifier based on Euclidean distance. (ii) DTW-1NN-I: It applies the nearest neighbor classifier based on dynamic time warping (DTW) that processes each dimension independently. and (iii) DTW-1NN-D: It applies the nearest neighbor classifier based on DTW that processes all dimensions simultaneously.
F.2 More detail about the Architecture
We use AdamW optimizer (Loshchilov & Hutter, 2017) with a fixed learning rate 1e-3 and a 1e-4 weight decay. We also use Lookahead scheduler (Zhang et al., 2019). Batch sizes are tuned based on the datasets since there are large differences in the dimension and length of each dataset.
As discussed in Section 3.3, we use Nyström Self-attention (Xiong et al., 2021) for accelerating the computation. Specifically, we set the embedding dimension $d_{model}=512$, the number of MHSA heads to $8$. For the Nystrom-based matrix approximation, we set the number of landmark points to 256 and the number of moore-penrose iterations for approximating pseudo-inverse to 6, which is recommended by its original paper. The final classifier consists of two fully connected layers: $\mathbb{R}^{L}\rightarrow\mathbb{R}^{L}\rightarrow\mathbb{R}^{C}$, where $C$ denotes the number of classes. We only feed the class token to the classifier. To facilitate the assumption of TimeMIL that treats MTSC as several one-vs-rest (OvR) binary classifications in the context of MIL, we use the binary cross entropy with one-hot encoding for the sequence label. We also adopt window-based random masking augmentation and a warm-up technique, as discussed below.
F.3 Window-based Random Masking Augmentation
This augmentation is only applied to the raw data in the training phase and aims to lead the model to learn the occlusion-invariant features. Recall the length of the input sequence is $T$. We first generate 10 non-overlapping windows with a size of $T/10$ that can fully cover the entire sequence. Suppose their indices are $[10]=\{1,2,\cdots,10\}$. For each iteration, we sample a set of windows $\mathcal{S}$ with a cardinality $|\mathcal{S}|=10p,p\in(0,1)$ from $[10]$ without replacement. Then, we set the time points covered by the windows in $\mathcal{S}$ to a random noise $\mathcal{N}(0,1)$. The example is shown in Fig. 8. We set $p\in\{0,0.5\}$ in our experiment.
Figure 8: The illustration of the window-based random masking. Top: The raw data of the input time sequence. Bottom: An example of the masking, where windows $1$ and $10$ is selected. The masked time points are marked in black.
F.4 warm-up Training Strategy
Considering the Transformer-like architecture learns slowly at the beginning, to facilitate the easy gradient flow to the backbone, we apply the following warm-up strategy. Recall the embedding of features is denoted as $\hat{\boldsymbol{x}}_{i}$, and the class token $\boldsymbol{x}_{i}^{\operatorname{cls}}$ after applying two transformer blocks. At the first few epochs (empirically set 10), we use $\alpha\mathbb{E}_{t}(\hat{\boldsymbol{x}}_{i}^{t})+(1-\alpha)\boldsymbol{x}_{i}^{\operatorname{cls}}$ to feed the final classifier, where $\alpha=0.99$. Afterwards, we still use $\boldsymbol{x}_{i}^{\operatorname{cls}}$ for the final classification.
Appendix G Additional Experiments
The additional results on all 30 UEA datasets are presented in Table 5. It is observed that classification-specific methods (TapNet, MOS-CNN, TodyNet, and Ours) often perform better. We also include Mamba (Gu & Dao, 2023), the recent state-of-the-art Selective State Spaces model. We implement the vanilla version of mamba with different numbers of blocks {1,2,4,8,10,22}, and the results are presented in Table 6.
Table 5: Comparison of different methods on 30 UEA datasets. (-) indicates the method is not able to obtain results due to memory limitations (we use an A100 40Gb GPU) on the EigenWorms dataset (dimension of 6 and length of 17894) and MotorImagery dataset (dimension of 64 and length of 3000).
Method Accuracy F1 Precision Recall AUC-ROC
Crossformer(-) 0.69 0.619 0.633 0.667 0.834
PatchTST 0.702 0.664 0.694 0.686 0.854
TimesNet 0.744 0.7 0.727 0.728 0.864
Dlinear 0.695 0.671 0.678 0.685 0.859
FEDformer 0.701 0.664 0.697 0.682 0.867
TapNet 0.759 0.745 0.764 0.751 0.832
MOS-CNN 0.78 0.764 0.788 0.765 0.863
TodyNet 0.762 0.744 0.759 0.751 0.869
Mamba-8 0.733 0.715 0.743 0.723 0.835
Ours 0.798 0.782 0.798 0.782 0.883
Table 6: Classification performance by Mamba with different number of blocks.
Method Accuracy F1 Precision Recall AUC-ROC
Mamba-1 0.725 0.704 0.729 0.713 0.827
Mamba-2 0.726 0.705 0.729 0.715 0.83
Mamba-4 0.729 0.711 0.73 0.718 0.832
Mamba-8 0.733 0.715 0.743 0.723 0.835
Mamba-10 0.727 0.705 0.727 0.714 0.833
Mamba-14 0.727 0.706 0.733 0.715 0.829
Mamba-22 0.726 0.708 0.727 0.716 0.829
Ours 0.798 0.782 0.798 0.782 0.883
Appendix H Synthetic Dataset
H.1 Dataset Generation
We simulate a binary dataset similar to noisy pulse signals. Consider the length of a sequence is 120. Again, $\boldsymbol{x}_{i}^{t}$ and ${y}_{i}^{t}$ denote the $t$th time point of the sequence $\boldsymbol{X}_{i}$ and its instance-level label. A negative sequence is generated as,
$\boldsymbol{x}_{i}^{t}\sim\mathcal{N}(0,0.5)\text{ and }{y}_{i}^{t}=0.$ (27)
A positive sequence, which consists of a noisy pulse, is generated as,
$\boldsymbol{x}_{i}^{t}\sim\left\{\begin{array}[]{l}\mathcal{N}(5,0.5)\text{ and }{y}_{i}^{t}=1\text{, for }t\in[a,a+20]\\ \mathcal{N}(0,0.5)\text{ and }{y}_{i}^{t}=0,\text{ otherwise}\end{array}\right.,$ (30)
where, $a\sim\mathcal{U}(55,65)$, is a random starting point for the pulse signal.
H.2 Decision Boundary
We randomly choose a positive sequence (i.e., a bag) and a negative one from the synthetic dataset. After applying feature extractor, both the positive sequence and negative sequence are projected onto a fixed-length (i.e., $L=128$) feature vectors (a positive bag $\boldsymbol{X}_{p}\in\mathbb{R}^{T\times 128}$ and a negative one $\boldsymbol{X}_{n}\in\mathbb{R}^{T\times 128}$). In this case, we have a total of 240 time points/instances ($T=240$). Subsequently, we apply PCA for these 240 instances, which reduces their dimensions from 128 to 2 for visualization, meaning we only use the first two principle components.

### v2

@@
 Contributions: (i) To the best of our knowledge, we are the first to formally formulate a generic MIL framework for multivariate and multi-class time series classification from an information-theoretic perspective.
-(ii) We propose a Time-aware MIL pooling based on a tokenized transformer and a novel learnable wavelet positional encoding to model complex patterns within time series.
+(ii) We propose a Time-aware MIL pooling based on a tokenized transformer and a novel learnable wavelet positional encoding (WPE) to model complex patterns within time series.
 The proposed method outperforms 26 recent state-of-the-art methods in 28 datasets and offers inherent interpretability.
@@
 Multivariate Time Series Classification.
-The recent DL methods specifically designed for MTSC can roughly be divided into two categories: (i) CNN/LSTM Hybrid architecture (Karim et al., 2019; Zhang et al., 2020), where LSTM is often used to capture sequential dependencies and CNN is used to capture the inter-variate correlations.
+The recent DL methods specifically designed for MTSC can roughly be divided into two categories: (i) CNN/LSTM Hybrid architecture (Karim et al., 2019; Zhang et al., 2020), where LSTM is often used to capture sequential dependencies and CNN is used to capture the local features.
 (ii) Purely CNN architecture (Ismail Fawaz et al., 2020; Li et al., 2021b; Tang et al., 2022), where long-term dependencies, short-term dependencies, and cross-channel dependencies are claimed to be captured by multiple kernels with varying kernel sizes.
-Recently, the authors of (Wu et al., 2023; donghao & wang xue, 2024) offered a model called General Time Series Analysis Framework, to be used in multiple mainstream tasks, including classification, imputation, short-term forecasting, long-term forecasting, and anomaly detection, with simple modifications.
+Recently, General Time Series Analysis Framework (Wu et al., 2023; donghao & wang xue, 2024) has been also proposed for multiple mainstream tasks, including classification, imputation, short-term forecasting, long-term forecasting, and anomaly detection, with simple modifications.
 Transformer-based models (Zhou et al., 2022; Zhang & Yan, 2023; Nie et al., 2023) and MLP-based models (Zeng et al., 2023; Zhang et al., 2022b; Li et al., 2023c, b) have also been developed and improved over the last few years for this purpose due to their excellent scaling behaviors.
@@
 In this section, we introduce three key components for applying the proposed TimeMIL to MTSC.
-First, we formulate the MTSC as a MIL problem in section 3.1 and 3.2.
-Second, we introduce a time-aware MIL pooling to capture the temporal ordering in time series through a wavelet-augmented transformer (Section 3.3).
+First, we formulate the MTSC as a MIL problem in Sec.
+3.1 and 3.2.
+Second, we introduce a time-aware MIL pooling to capture the temporal ordering in time series through a wavelet-augmented transformer (Sec.
+3.3).
+Third, we introduce how to quantify the importance of instances in Sec.
+3.4.
 The entire framework of the proposed TimeMIL is depicted in Fig.
 2.
-Third, we introduce interpretable pooling to assess the importance of instance, as discussed in section 3.4.
 A summary of the proposed framework is presented in Algorithm 1.
 3.1 Problem Formulation
-Multivariate Time Series data is typically presented as $\{\boldsymbol{X}_{1},\cdots,\boldsymbol{X}_{n}\}$, where $\boldsymbol{X}_{i}=\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}$ is a time series with $T$ time points, with each time point $\boldsymbol{x}_{i}^{t}\in\mathbb{R}^{d}$ being a $d$-dimensional vector.
+Multivariate time series data is typically presented as $\{\boldsymbol{X}_{1},\cdots,\boldsymbol{X}_{n}\}$, where $\boldsymbol{X}_{i}=\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}$ is a time series contains $T$ time points, with each time point $\boldsymbol{x}_{i}^{t}\in\mathbb{R}^{d}$ being a $d$-dimensional vector.
 It is noteworthy that the time points are shift-variant and ordered.
-Similar to the standard classification task, the goal of MTSC is to learn a direct mapping from feature space $\mathcal{X}$ to label space $\mathcal{Y}$ using the training data set $\{(\boldsymbol{X}_{1},y_{1}),\cdots,(\boldsymbol{X}_{n},y_{n})\}$, where $y_{i}$ is the label for a single time series.
+The goal of MTSC is to learn a direct mapping from feature space $\mathcal{X}$ to label space $\mathcal{Y}$ using the training data $\{(\boldsymbol{X}_{1},y_{1}),\cdots,(\boldsymbol{X}_{n},y_{n})\}$, where $y_{i}$ is the label for each time series.
 3.2 MTSC as A MIL Problem
@@
 This hinders the direct translation of classic MIL into MTSC tasks.
-To address this limitation, we propose a time-aware MIL pooling in Section 3.3.
+To address this limitation, we propose a time-aware MIL pooling in Sec.
+3.3.
+Algorithm 1 Time-aware MIL (Forward Propagation)
+0:  Input sequence $\boldsymbol{X}_{i}=\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}$.
+A neural network contains: $f$: feature extractor, $g$: bag-level classifier, $\operatorname{WPE}$ bases, and class token $\boldsymbol{x}^{\operatorname{cls}}$.
+0:  Predicted label $\hat{y}_{i}$.
+1:  Get embedding: $\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}\leftarrow f(\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\})$MIL pooling from here (Section 3.3)
+2:  Init class token for the bag: $\boldsymbol{x}_{i}^{\operatorname{cls}}\leftarrow\boldsymbol{x}^{\operatorname{cls}}$.
+3:  Add class token: $\boldsymbol{X}_{i}^{\operatorname{cls}}=\boldsymbol{x}_{i}^{\operatorname{cls}}\cup\boldsymbol{X}_{i}$.
+4:  for j=1:2 do
+5:     PE: $\boldsymbol{X}_{i}\leftarrow\boldsymbol{X}_{i}+WPE_{j}(\boldsymbol{X}_{i})$.
+# $WPE_{j}(\boldsymbol{X}_{i})$ is from Eq.
+10.
+Here, $\boldsymbol{X}_{i}$ is the part of $\boldsymbol{X}_{i}^{\operatorname{cls}}$.
+6:     $\boldsymbol{X}_{i}^{\operatorname{cls}}\leftarrow Transformer_{j}(\boldsymbol{X}_{i}^{\operatorname{cls}})$.
+7:  end for#Bag-level classification from here
+8:  $\hat{y}_{i}\leftarrow g(\boldsymbol{x}_{i}^{\operatorname{cls}})$. #$\boldsymbol{x}_{i}^{\operatorname{cls}}$ is obtained from $\boldsymbol{X}_{i}^{\operatorname{cls}}$.
 3.3 Time-Aware MIL Pooling for MTSC
 From an entropy perspective, understanding permutation-variant properties in time series can be quite insightful.
-As discussed in Section 3.2, we assume that each time point (instance) $\boldsymbol{x}_{i}^{j}$ in a time series (bag) is a realization of a random variable $\Theta^{j}$ conditioned by a time index $t^{j}$.
+As discussed in Sec.
+3.2, we assume that each time point (instance) $\boldsymbol{x}_{i}^{j}$ in a time series (bag) is a realization of a random variable $\Theta^{j}$ conditioned by a time index $t^{j}$.
 The resulting bag can be represented as a random variable $\boldsymbol{X}=\{(\Theta^{1}|t^{1}),\cdots,(\Theta^{T}|t^{T})\}$, where $t^{j}\neq t^{k},\forall j\neq k$.
@@
 Specifically, we use the approximation of self-attention proposed by (Xiong et al., 2021) in our implementation to reduce the complexity of the proposed TimeMIL.
-Figure 4: The proposed learnable wavelet positional encoding: First, wavelet transform is performed for the input signal (by excluding the class token) with each wavelet basis (Eq.
-17).
-Second, the signals are aggregated in the wavelet domain by a summation (Eq.
-13).
-In the case of $n_{w}=3$, we use 3 learnable wavelet bases ($\boldsymbol{\Psi}_{1},\boldsymbol{\Psi}_{2},\boldsymbol{\Psi}_{3}$) to model changing frequency and time scales.
 Wavelet positional encoding.
@@
 $\Phi(\tilde{\boldsymbol{X}}_{i},\boldsymbol{\Psi}_{j})=\left[\begin{array}[]{c}\tilde{\boldsymbol{X}}_{i1}\circledast\psi_{a_{j1},b_{j1}}\\ \vdots\\ \tilde{\boldsymbol{X}}_{iL}\circledast\psi_{a_{jL},b_{jL}}\end{array}\right]^{\top}\in\mathbb{R}^{n\times L}.$ (17)
+The resulting WPE is depicted in Fig. 4.
 3.4 Interpretability
 The proposed time-aware MIL is naturally interpretable due to its ability to localize time points of interest within a time series.
-This is achieved by hacking into the attention map of the transformer layers in the proposed time-aware MIL pooling.
+One way to achieve this is to hack into the attention map of the transformer layers in the proposed time-aware MIL pooling.
 The attention map (refer to Eq.
@@
 We only use the class token to calculate the importance weight, as it determines the most relevant time points that characterize a certain class label in MTSC tasks.
-Algorithm 1 Time-aware MIL (Forward Propagation)
-0:  Input sequence $\boldsymbol{X}_{i}=\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}$.
-A neural network contains: $f$: feature extractor, $g$: bag-level classifier, $WPE$ bases, and $\boldsymbol{x}^{\operatorname{cls}}$: class token.
-Two Transfomer blocks.
-0:  Predicted label $\hat{y}_{i}$.
-1:  Get embedding: $\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}\leftarrow f(\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\})$MIL pooling from here (Section 3.3)
-2:  Init class token for the bag: $\boldsymbol{x}_{i}^{\operatorname{cls}}\leftarrow\boldsymbol{x}^{\operatorname{cls}}$.
-3:  Add class token: $\boldsymbol{X}_{i}^{\operatorname{cls}}=\boldsymbol{x}_{i}^{\operatorname{cls}}\cup\boldsymbol{X}_{i}$.
-4:  for j=1:2 do
-5:     PE: $\boldsymbol{X}_{i}\leftarrow\boldsymbol{X}_{i}+WPE_{j}(\boldsymbol{X}_{i})$.
-# $WPE_{j}(\boldsymbol{X}_{i})$ is from Eq.
-10.
-Here, $\boldsymbol{X}_{i}$ is the part of $\boldsymbol{X}_{i}^{\operatorname{cls}}$.
-6:     $\boldsymbol{X}_{i}^{\operatorname{cls}}\leftarrow Transformer_{j}(\boldsymbol{X}_{i}^{\operatorname{cls}})$.
-7:  end for#Bag-level classification from here
-8:  $\hat{y}_{i}\leftarrow g(\boldsymbol{x}_{i}^{\operatorname{cls}})$. #$\boldsymbol{x}_{i}^{\operatorname{cls}}$ is obtained from $\boldsymbol{X}_{i}^{\operatorname{cls}}$.
+Figure 4: The proposed learnable wavelet positional encoding: First, wavelet transform is performed for the input signal (by excluding the class token) with each wavelet basis (Eq.
+17).
+Second, the signals are aggregated in the wavelet domain by a summation (Eq.
+13).
+In the case of $n_{w}=3$, we use 3 learnable wavelet bases ($\boldsymbol{\Psi}_{1},\boldsymbol{\Psi}_{2},\boldsymbol{\Psi}_{3}$) to model changing frequency and time scales.
 Table 1: Results of Group 1 Experiments.
@@
 4 Experiments
-4.1 Setup and Baselines
+4.1 Experimental setup and Baselines
 We use the UEA benchmark datasets to validate the superiority of the proposed TimeMIL.
@@
 The proposed method demonstrates superiority over other recently proposed competing methods in both Group 1 and Group 2 experiments (refer to Table 1 and 2).
+Please refer to Appendix G for additional results.
+Figure 5: Exemplary attention maps learned in TimeMIL using different datasets (rows) including synthetic dataset, StandWalkJump dataset, and AtrialFibrillation dataset, featuring distinct patterns of interest (columns).
+TimeMIL accurately localized patterns of interest.
 Group 1 results.
@@
 Specifically, our method achieved a 77.7% average accuracy, surpassing the other methods in 5 out of 10 datasets and achieving an average rank of 2.85 out of a total of 17 methods.
-Notably, the proposed method outperforms the recent state-of-the-art ModernTCN (donghao & wang xue, 2024) by 3.5% in average accuracy and 1.15 in average rank.
+Remarkably, the proposed method outperforms the recent state-of-the-art ModernTCN (donghao & wang xue, 2024) by 3.5% in average accuracy and 1.15 in average rank.
 Table 3: Comparison of different MIL pooling (Left) and positional encoding (Right) with Group 1 experiments.
@@
 WPE (Ours) 0.745 0.774
-Figure 5: Exemplary attention maps learned in TimeMIL using different datasets (rows) including synthetic dataset, StandWalkJump dataset, and AtrialFibrillation dataset, featuring distinct patterns of interest (columns).
-TimeMIL accurately localized patterns of interest.
-Figure 6: Decision boundary learned in fully supervised method (Left) versus TimeMIL (Right) using the synthetic dataset.
 4.3 Ablation on Model Design Variants
@@
 5 and 6.
+Figure 6: Decision boundary learned in fully supervised method (Left) versus TimeMIL (Right) using the synthetic dataset.
 Decision boundary.
@@
 The instance-level decision-making mechanism makes the proposed TimeMIL inherently interpretable.
-Leveraging this, we can obtain the importance score (Eq.
-18).
+Leveraging this, we can obtain the importance score.
 TimeMIL accurately localized patterns of interest (i.e., positive instances) in the UEA dataset (Fig.
@@
 As a result, TimeMIL demonstrates superiority over 26 methods in 28 datasets and illustrates impressive interpretability by accurately localizing patterns of interest.
-6 Limitation
-Currently, our work primarily focuses on modeling the sparsity and locality of time points, and we have not yet extended the interpretability pooling to the channel level.
-Our work may also be limited by the case that a time point fuses multiple variates that represent potential delayed events or measurements.
-We envision that a major difference in expanding to the channel level would involve designing a channel-based positional encoding.
-This will be a topic of exploration in our future work.
-7 Impact Statements
+Impact Statements
 This paper presents a novel TimeMIL framework to model the temporal correlation and ordering, leveraging a tokenized transformer with a specialized learnable wavelet positional token.
-The potential societal consequences of TimeMIL can be summarised three-fold: (i) Theoretical View: Commencing with a viewpoint in weakly supervised learning, we are the first to formally introduce MIL into time series-related tasks.
+The potential societal consequences of TimeMIL can be summarized three-fold:
+(i) Theoretical View.
+Commencing with a viewpoint in weakly supervised learning, we are the first to formally introduce MIL into time series-related tasks.
 We rigorously examine feasibility from an information-theoretic perspective.
-In response to the various limitations of MIL in MTSC, we propose corresponding modules to preserve the intrinsic temporal correlation and ordering properties within time series.
-We draw connections between the MIL and MTSC.
-This provides an improved justification, relating it to existing MTSC methods.
-(ii) Flexibility: Our framework significantly expands the MTSC applications.
-For example, future work involving financial forecasting, predictive maintenance, anomaly detection, healthcare monitoring, financial forecasting, environmental monitoring, and speech and signal processing, among others, can all be incorporated within this framework.
-(iii) Interpretability: Compared with the previously existing methods, TimeMIL could provide the interpretable interest of the pattern of the network.
+To address a variety of limitations in applying MIL in MTSC, we propose a time-aware MIL pooling to preserve the intrinsic temporal correlation and ordering properties within time series.
+In summary, we draw theoretical connections between the MIL and MTSC.
+(ii) Applicability.
+Our framework potentially expands the applications of MTSC to financial forecasting, predictive maintenance, anomaly detection, healthcare monitoring, financial forecasting, environmental monitoring, and speech and signal processing.
+(iii) Interpretability.
+Compared with the previously existing methods, TimeMIL could provide the interpretable interest of the pattern of the network.
 By visualizing the points of time series that strongly influence predictions, TimeMIL can assist in ensuring model robustness.
-This can help in identifying vulnerabilities and reducing the risk of adversarial attacks.
+This can help identify vulnerabilities and reduce the risk of adversarial attacks.
 Meanwhile, interpretability contributes to the explainability of MTSC, making it easier for researchers, practitioners, and stakeholders to comprehend/validate how and why a model makes specific predictions.
+Limitation.
+We have not yet extended the MIL to consider the cross-channel information.
+We envision a major difference in expanding the current framework to the multi-channel version of TimeMIL, which could involve designing cross-channel temporal attention and positional encoding.
+This will be a topic of exploration in future work.
+Acknowledgement
+This material is based upon the work supported by the National Science Foundation under Grant Number CNS-2204721.
+It is also supported by our collaborative project with MIT Lincoln Lab under Grant Numbers 2015887 and 7000612889.
 Appendix A UEA Datasets Detail
@@
 We also adopt window-based random masking augmentation and a warm-up technique, as discussed below.
+The importance score can be conveniently approximated by using Average-Pooling Based Attention (APBA) proposed by (Zhu et al., 2023).
 F.3 Window-based Random Masking Augmentation
@@
 We also include Mamba (Gu & Dao, 2023), the recent state-of-the-art Selective State Spaces model.
-We implement the vanilla version of mamba with different numbers of blocks {1,2,4,8,10,22}, and the results are presented in Table 6.
+We implement the vanilla version of mamba with different numbers of blocks {1,2,4,8,10,14,22}, and the results are presented in Table 6.
 Table 5: Comparison of different methods on 30 UEA datasets.
@@
 Mamba-8 0.733 0.715 0.743 0.723 0.835
-Ours 0.798 0.782 0.798 0.782 0.883
+Ours 0.791 0.782 0.790 0.782 0.883
 Table 6: Classification performance by Mamba with different number of blocks.
@@
 Mamba-22 0.726 0.708 0.727 0.716 0.829
-Ours 0.798 0.782 0.798 0.782 0.883
+Ours 0.791 0.782 0.790 0.782 0.883
 Appendix H Synthetic Dataset